# Focus On This, Not That! Steering LLMs With Adaptive Feature Specification

### v2

@@ A.3 Evaluation Metrics
Computing the focus accuracy metric. We report the accuracy of generations when evaluating FT models. As we are guaranteed to include the task labels within the model’s response through constrained decoding, we simply check to see if the focus label, $y_{\text{focus}}$, is within the model’s response or not in order to determine if the model’s response is correct.
A.4 FIT Focus Instructions and Prompt Templates
Focus instructions $\text{focus}(F_{i})$:
Prompt template for SS. We consider the following prompt templates for the SS dataset:
@@ A.4 FIT Focus Instructions and Prompt Templates
where for the instruction with no focus instruction, $\emptyset$, we simply drop the feature considerations section and focus instruction from the prompt.
Prompt template for SMNLI. We consider the following prompt templates for the SMNLI dataset.
@@
where we again drop the feature considerations section and focus instruction from the prompt to form the instruction prompt $\emptyset$ with no focus instruction.
BBQ focus instruction prompt template $\mathcal{I}_{\text{focus}}$ We consider the following prompt templates for the BBQ dataset.
@@
where we again drop the feature considerations to get the template for the focus instruction $\emptyset$.
A.5 Baselines
We include results for the following two baselines to fruther supplement the results presented in the main section of the paper.
@@ A.5 Baselines
For completeness, we report two additional baselines: vanilla SFT and zero-shot baselines.
@@
A.6 Additional Baselines Results
In this section, we include the two additional baselines -SFT($y$) and zero-shot - on the SMNLI dataset to further supplement the results in Section 4.2.
@@ A.6 Additional Baselines Results
A.7 SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
We analyse the impact of using the same versus different sets of focus instructions at training and test time when applying FIT models. Specifically, we generate alternative test set focus instructions by paraphrasing the training focus instructions, as shown in Section A.4, using ChatGPT.
@@ A.7 SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
A.8 Comparison of FIT against a specific debiasing technique
@@
Experiment. To explore FIT’s broader applicability, we compare its performance as a debiasing method against a well-known debiasing technique: the Product of Experts (PoE) method (Mahabadi et al., 2020). PoE involves training a bias model $f_{B}$, which is trained exclusively on bias features. This bias model mediates the training of the final model $f$ by combining their predictions through an elementwise product: $\sigma(f(x))\odot\sigma(f_{B}(x_{B}))$, where $x\in\mathcal{D}$ and $x_{B}$ represents the biased feature of $x$.
We adapted this approach to our setting by training a bias model on the stereotypical labels within the BBQ dataset. These labels correspond to group-stereotypical associations. For autoregressive models, we further modified the PoE method by extracting and normalizing the logits of the first newly generated token position over the set of single tokens representing the answer options.
Results. The results of the debiasing experiment comparing FIT to the PoE method is shown in Figure 11. FIT performs equally as well as the PoE method as shown by the comparing the default prompt accuracy () for the PoE models against the $\text{focus}(C)$ results for the FIT models; both metrics correspond to causal accuracy for these prompt types. Indicating that FIT performs just as well as a dedicated debiasing technique.
@@
A.9 Spurious Sentiment (SS) Dataset
@@
Finally, we prove that $\rho_{\text{spurious}}$ gives the coocurrence rate/predictivity between the label $Y$ and the spurious feature $S$, and is well-defined notation in the sense that it corresponds to Definition 1 so that $\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$, where $y_{s}$ is the label that spurious feature value $s$ is spuriously correlated with.
Proposition 1.
@@
Within our experiments in Section 4.1, we always force the label distribution to be balanced, that is $\rho_{C}=1/2$, and assume that within each dataset split, $\rho_{\text{spurious}}$ is the same rate for all spurious feature values.
Independence conditions during training for FIT. As specified in Section A.1, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features. Here, $Y_{S}$ is the spurious label spuriously correlated to spurious feature value $S$. The results below give sufficient conditions for these independence conditions to be specified with respect to the DGP described in Figure 12.
Proposition 2.
@@
A.10 Spurious NLI dataset (SMNLI)
We generate a tertiary NLI dataset, SMNLI, with a known spurious feature. We do this considering the MNLI dataset Williams et al. (2018). This is a NLI dataset with three labels: entailment (0), neutral (1) and contradiction (2), where data is sampled from 5 underlying categories or genres (telephone, government, travel, fiction or slate). We aim to induce spurious correlations between the underlying genres and labels.
@@ A.10 Spurious NLI dataset (SMNLI)
We show that the notion of $\rho_{\text{spurious}}$ in Equation 38 aligns with the notation in Definition 1.
Proposition 4.
@@
Independence conditions during training for FIT. As specified in Section A.1, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features, where, again, $Y_{S}$ is the label spuriously associated to spurious feature value $S$. The results below give sufficient conditions for this to occur with respect to the DGP described in Figure 14.
Proposition 5.
@@
From Proposition 4, we have that $\mathbb{P}(Y=y_{s}\mid S=s)=1/3$ when $\rho_{\text{spurious}}=1/3$ for any spurious feature value $s\in\text{Val}(S)$. Moreover, we have that from Equation 39
@@
$\mathbb{P}(C=c)$ $=\sum_{s}\underbrace{\mathbb{P}(C=c\mid S=s)}_{=\frac{1}{3}}\underbrace{\mathbb{P}(S=s)}_{\frac{1}{k}}$ (48)
@@
A.11 Spurious HANS dataset (SHANS)
We generate a binary NLI dataset, SHANS, with a known spurious feature. We do this considering the HANS dataset McCoy (2019). This is an NLI data set with two labels: entailment (0) and contradiction (1). This is an adversarial dataset designed to assess different NLI models’ reliance on spurious heuristics rather than on the underlying relationship between the premise and the hypothesis when making predictions. Specifically, the author’s consider three major categories of heuristics: lexical overlap heuristic (assuming that a premise entails from words within the hypothesis) , sub-sequence heuristic( assuming that the premise entails all any of its contiguous sub-sequences of words) and constituent heuristic (assuming that a premise entails a hypothesis that is any constituent within it’s syntactic parse tree).
@@ A.11 Spurious HANS dataset (SHANS)
Transferred results from SMNLI. As we use the same DGP as for the SMNLI dataset described in Section A.11, all of the results that we have proven for SMNLI, translate to the SHANS dataset. In particular, we have that $\rho_{\text{spurious}}$ aligns with the notation in Definition 1, and that we have $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ under the assumptions of balanced label distributions and $\rho_{\text{spurious}}=1/2$ within the training set used for FIT training.
@@ A.12 FIT on SHANS
Spurious HANS (SHANS) dataset. We generate binary NLI dataset sub-sampled from HANS (McCoy, 2019), a dataset designed to challenge NLI models by exposing common heuristics they rely on, such as lexical overlap (whether the hypothesis shares many words with the premise), sub-sequence (whether the hypothesis is a contiguous sub-sequence of the premise), and constituent (whether the hypothesis is a grammatical sub-structure of the premise). The presence of these heuristics are spuriously correlated with labels through sub-sampling the presence of each of the heuristics from the original dataset. The degree of co-occurrence is governed by $\rho_{\text{spurious}}$, which varies according to the test sets described in Section 3. We ensure that $\rho_{\text{spurious}}$ is the same for all feature values within each dataset split. In particular, we set $\rho_{\text{spurious}}$ to be $0.5,0.5,0.9,0.25$ and $0.9$ on $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{iid}}$, $\mathcal{D}_{\text{high}}$, $\mathcal{D}_{\text{low}}$ and $\mathcal{D}_{\text{flipped}}$ respectively.

### v3

@@ Appendix D Evaluation Metrics
Computing the focus accuracy metric. We report the focus accuracy $\mathcal{A}_{\text{focus}}$ of generations when evaluating FT models. As we are guaranteed to include the task labels within the model’s response through constrained decoding, we simply check to see if the focus label, $y_{\text{focus}}$, is within the model’s response or not in order to determine if the model’s response is correct.
Appendix E FIT Focus Instructions and Prompt Templates
Prompt templates. Here we provide the prompt templates that we use for FIT. Note, that we drop the feature considerations section and focus instruction from the prompt to form the instruction prompt $\emptyset$ with no focus instruction.
@@ Appendix E FIT Focus Instructions and Prompt Templates
Focus instructions $\text{focus}(F_{i})$:
Appendix F Additional Baselines Results
In addition to the baselines that we present in the main paper, Few-shot and $\text{SFT}(y_{\text{focus}})$, we include two additional baselines to further supplement these results. We give the complete list of baselines that we consider below:
@@ Appendix F Additional Baselines Results
We give the full set of results for all datasets and models across the complete set of baselines listed above in Figure 6, Figure 7, and Figure 8.
@@
Appendix G SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
We analyse the impact of using the same versus different sets of focus instructions at training and test time when applying FIT models. Specifically, we generate alternative test set focus instructions by paraphrasing the training focus instructions, as shown in Appendix E, using ChatGPT.
@@ Appendix G SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
Appendix H Comparison of FIT against a specific debiasing technique
@@
Experiment. To explore FIT’s broader applicability, we compare its performance as a debiasing method against a well-known debiasing technique: the Product of Experts (PoE) method (Mahabadi et al., 2020). PoE involves training a bias model $f_{B}$, which is trained exclusively on bias features. This bias model mediates the training of the final model $f$ by combining their predictions through an elementwise product: $\sigma(f(x))\odot\sigma(f_{B}(x_{B}))$, where $x\in\mathcal{D}$, for dataset $\mathcal{D}$, and $x_{B}$ represents the biased feature of $x$.
We adapted this approach to our setting by training a bias model on the stereotypical labels within the BBQ dataset. These labels correspond to group-stereotypical associations. For autoregressive models, we further modified the PoE method by extracting and normalising the logits of the first newly generated token position over the set of single tokens representing the answer options.
Results. The results of the debiasing experiment comparing FIT to the PoE method is shown in Figure 10. FIT performs equally as well as the PoE method as shown by the comparing the default prompt accuracy ($\emptyset$) for the PoE models against the $\text{focus}(C)$ results for the FIT models; both metrics correspond to causal accuracy for these prompt types. Indicating that FIT performs just as well as a dedicated debiasing technique.
@@
Appendix I Spurious Sentiment (SS) Dataset
@@
Finally, we prove that $\rho_{\text{spurious}}$ gives the coocurrence rate/predictivity between the label $Y$ and the spurious feature $S$, and is well-defined notation in the sense that it corresponds to Definition 3.1 so that $\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$, where $y_{s}$ is the label that spurious feature value $s$ is spuriously correlated with.
Proposition I.1.
@@
Within our experiments in Section 4.1, we always force the label distribution to be balanced, that is $\rho_{C}=1/2$, and assume that within each dataset split, $\rho_{\text{spurious}}$ is the same rate for all spurious feature values. Independence conditions during training for FIT. As specified in Appendix C, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features. Here, $Y_{S}$ is the spurious label spuriously correlated to spurious feature value $S$. The results below give sufficient conditions for these independence conditions to be specified with respect to the DGP described in Figure 11.
Proposition I.2.
@@
Appendix J Spurious NLI dataset (SMNLI)
We generate a tertiary NLI dataset, SMNLI, with a known spurious feature. We do this considering the MNLI dataset (Williams et al., 2018). This is a NLI dataset with three labels: entailment (0), neutral (1) and contradiction (2), where data is sampled from 5 underlying categories or genres (telephone, government, travel, fiction or slate). We aim to induce spurious correlations between the underlying genres and labels.
@@
We show that the notion of $\rho_{\text{spurious}}$ in Equation 43 aligns with the notation in Definition 3.1.
Proposition J.1.
@@
Independence conditions during training for FIT. As specified in Appendix C, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features, where, again, $Y_{S}$ is the label spuriously associated to spurious feature value $S$. The results below give sufficient conditions for this to occur with respect to the DGP described in Figure 13.
Proposition J.2.
@@
From Proposition J.1, we have that $\mathbb{P}(Y=y_{s}\mid S=s)=1/3$ when $\rho_{\text{spurious}}=1/3$ for any spurious feature value $s\in\text{Val}(S)$. Moreover, we have that from Equation 44
@@
$\mathbb{P}(C=c)$ $=\sum_{s}\underbrace{\mathbb{P}(C=c\mid S=s)}_{=1/3}\underbrace{\mathbb{P}(S=s)}_{1/k}$ (53)
@@
Appendix K Spurious HANS dataset (SHANS)
We generate a binary NLI dataset, SHANS, with a known spurious feature. We do this considering the HANS dataset (McCoy, 2019). This is an NLI data set with two labels: entailment (0) and contradiction (1). This is an adversarial dataset designed to assess different NLI models’ reliance on spurious heuristics rather than on the underlying relationship between the premise and the hypothesis when making predictions. Specifically, the author’s consider three major categories of heuristics: lexical overlap heuristic (assuming that a premise entails from words within the hypothesis) , sub-sequence heuristic( assuming that the premise entails all any of its contiguous sub-sequences of words) and constituent heuristic (assuming that a premise entails a hypothesis that is any constituent within it’s syntactic parse tree).
@@
Transferred results from SMNLI. As we use the same DGP as for the SMNLI dataset described in Appendix K, all of the results that we have proven for SMNLI, translate to the SHANS dataset. In particular, we have that $\rho_{\text{spurious}}$ aligns with the notation in Definition 3.1, and that we have $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ under the assumptions of balanced label distributions and $\rho_{\text{spurious}}=1/2$ within the training set used for FIT training.
@@
Spurious HANS (SHANS) dataset. We generate binary NLI dataset sub-sampled from HANS (McCoy, 2019), a dataset designed to challenge NLI models by exposing common heuristics they rely on, such as lexical overlap (whether the hypothesis shares many words with the premise), sub-sequence (whether the hypothesis is a contiguous sub-sequence of the premise), and constituent (whether the hypothesis is a grammatical sub-structure of the premise). The presence of these heuristics are spuriously correlated with labels through sub-sampling of the presence of each of the heuristics from the original dataset. The degree of co-occurrence is governed by $\rho_{\text{spurious}}$, which varies according to the test sets described in Section 3. We ensure that $\rho_{\text{spurious}}$ is the same for all feature values within each dataset split. In particular, we set $\rho_{\text{spurious}}$ to be $0.5,0.5,0.9,0.25$ and $0.9$ on $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{iid}}$, $\mathcal{D}_{\text{high}}$, $\mathcal{D}_{\text{low}}$ and $\mathcal{D}_{\text{flipped}}$ respectively.